A copper wire of radius 0.1 mm and resistance 1 kΩ is connected across a power supply of 20 V. (a) How many electrons are transferred per second between the supply and the wire at one end? (b) Write down the current density in the wire.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. r = 0.1 mm = 10^-4 R = 1 KΩ = 10^3 Ω, V = 20 V a) No.of electrons transferred i = V/R = 20/10^3 = 20 * 10^-3 = 2 * 10^-2 A q = i t = 2 * 10^-2 * 1 = 2 * 10^-2 C. No. of electron transferred = 2 * 10^-2/1.6 * 10^-19 = 2 * 10^-17/1.6 = 1.25 * 10^17. b)Current density of wire = i/A = 2 * 10^-2/π * 10^10^-8 = 2/3.14 * 10^+6 0.6369 * 10^+6 = 6.37 * 10^5 A/m^2.